# Transition to Petschek Reconnection in Subrelativistic Pair Plasmas: Implications for Particle Acceleration

### v2

@@ 1 Introduction
Reconnection is a ubiquitous plasma process that rapidly converts magnetic energy into particle energy via rearrangement of the magnetic field topology. Originally posited to explain solar flares, reconnection now plays a key role in the theories of a multitude of astrophysical systems, from the heliosphere to compact objects. Observations and simulations suggest that reconnection can energize particles far beyond thermal energies, generating hard power-law energy spectra (L. Sironi & A. Spitkovsky, 2014; F. Guo et al., 2014). The conditions under which this particle acceleration occurs, the precise mechanisms responsible, and the shape of the resultant spectra have been intensively studied in recent years with the aid of particle-in-cell (PIC) simulations (X. Li et al., 2021; F. Guo et al., 2024; L. Sironi et al., 2025). These simulations have enabled the study of kinetic reconnecting plasmas on increasingly large spatial and temporal scales.
@@
One of the key determinants of particle acceleration is the current sheet geometry: each particle acceleration mechanism requires certain configurations of electric and magnetic fields to operate. There are two broad categories of reconnecting current sheet geometries: single X-point and multiple X-point111also referred to as X-lines, the three-dimensional analogues of X-points. (H. Ji & W. Daughton, 2011). In the former, there is just one diffusion region where the field lines break and reconnect, while most of the reconnecting plasma proceeds directly from the inflow to the outflow exhaust. In the latter, the current sheet fragments due to tearing or plasmoid instability, resulting in a dynamic chain of magnetic islands with X-points in between (N. F. Loureiro et al., 2007). [^2] In this paper we shall use X-point and diffusion region interchangeably to refer to areas where non-ideal electric fields dissipate inflowing oppositely-aligned magnetic flux. Recent work has established the importance of direct electric field acceleration at X-points in relativistic plasmas as an injection mechanism for subsequent acceleration (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025), a pathway which is more accessible in systems with more X-points. Furthermore, plasmoid-based Fermi acceleration mechanisms at later stages of energization (J. F. Drake et al., 2006; F. Guo et al., 2019; H. Hakobyan et al., 2021) are also primarily operative in the multiple X-point regime. In contrast, pickup acceleration occurs mainly in the single X-point regime for particles that do not pass through the X-point (J. F. Drake et al., 2009).
@@
Given the strong dependence of particle acceleration on the current sheet geometry, it would be useful to have a framework to predict the transition from single X-point reconnection to multiple X-point reconnection. In the MHD model, this is accomplished by the Lundquist number $S\equiv Lv_{A}/(\eta c/4\pi)$, where $L$ is the half-length of the current sheet, $v_{A}$ is the Alfvén speed, and $\eta$ is the resistivity, assumed to be constant. When $S>S_{c}\sim 10^{4}$, the reconnecting current sheet fragments, resulting in fast plasmoid-mediated multiple X-point reconnection (H. Ji & W. Daughton, 2011). Although recent strides have been made in understanding the tearing properties of collisionless current sheets (M. Hoshino, 2020; K. M. Schoeffler et al., 2024; K. Akutagawa et al., 2025), there is no comparable metric to a priori predict the global reconnection geometry in collisionless systems, suggesting the need for kinetic studies.
@@
These questions also tie in closely to debates concerning the origin of fast reconnection. H. E. Petschek (1964) proposed that the slow Sweet-Parker reconnection (E. N. Parker, 1957) could be made fast by flanking a small diffusion region with four slow-mode shocks responsible for the majority of energy dissipation. His model has proven controversial since, in the absence of additional physics to localize the diffusion region, the Petschek solution is unstable and devolves to that of Sweet and Parker in MHD simulations (D. A. Uzdensky & R. M. Kulsrud, 2000; T. G. Forbes et al., 2013). On the other hand, slow-mode shocks have been observed in the Earth’s magnetotail for decades (e.g., W. C. Feldman et al., 1984; Y. Saito et al., 1995; N. K. Walia et al., 2024), providing observational evidence in favor of Petschek’s theory. Large-domain kinetic simulations would thus be well suited to address whether Petschek-type outflows can form, and, if so, under what conditions.
@@
where $B_{0}$ and $n_{0}$ are the magnetic field and number density of the upstream plasma, respectively. The magnetization represents the characteristic Lorentz factor that particles would gain if all the magnetic energy were dissipated into cold particles. Thus, $\sigma\gg 1$ indicates relativistic reconnection and $\sigma<1$ the subrelativistic regime. High-$\sigma$ systems are characterized by rapid plasmoid production, growth, and mergers (L. Sironi et al., 2016). Since the rate of plasmoid formation increases with increasing $\sigma$, by lowering $\sigma$ below 1 we can address systems with lower tearing rates than those typically studied for pair plasmas. The result, as we shall demonstrate, is that while high-$\sigma$ simulations always yield multiple X-point plasmoid chains, those with low-$\sigma$ instead have just a single X-point. By comparing otherwise identical simulations that differ only in $\sigma$, we can understand the bifurcation of the system into single or multiple X-point regimes and study how this bifurcation affects the resultant particle acceleration.
In this investigation, we focus on two magnetization values: $\sigma=0.3$ and $\sigma=10$, which have (upstream) Alfvén speeds $v_{A}=c\sqrt{\sigma/(\sigma+1)}$ of $\approx 0.48c$ and $\approx 0.95c$, respectively. The $\sigma=10$ case is well within the relativistic reconnection regime and, as we shall see, the resulting reconnection rate, current sheet structure, and particle energy spectra are consistent with the existing literature (e.g., L. Sironi et al., 2016; F. Guo et al., 2024). This acts as a control with which we may contrast the $\sigma=0.3$ case. Although we focus on only these two cases, we expect that the key results for $\sigma=10$ also apply to $\sigma>10$, and the results of $\sigma=0.3$ apply to $\sigma<0.3$, with the range $0.3<\sigma<10$ presenting an intermediate regime.
@@ 2 Simulation Setup
We perform 2D simulations of reconnection using the electromagnetic PIC code TRISTAN-V2 (H. Hakobyan et al., 2024). For all simulations, the mass ratio $m_{i}/m_{e}=1$, and there are 16 total particles per cell. The speed of light is set to $c=0.45$ cells per time step to satisfy the CFL condition. We vary the magnetization, skin depth, and domain size as described in Table 1.
@@ 3.1 Petschek-like Exhausts
Reducing the magnetization from $\sigma=10$ to $\sigma=0.3$ results in a stark change in the structure of the current sheet, as can be seen in Figure 1. In the $\sigma=10$ case, as expected, the tearing instability results in copious plasmoid formation, frequent plasmoid mergers, and rapid plasmoid growth. The result is a self-similar current sheet made up of plasmoids of various sizes, as described in detail in L. Sironi et al. (2016). This is a clear case of multiple X-point reconnection wherein field lines reconnect at many locations along the current sheet. However, a primary X-point can be identified at $y\approx 360\;c/\omega_{p}$ during this timestep.
@@
The exhausts are even more pronounced in Run V that has a domain size ten times longer in the $y$-direction, shown in Figure 2. Here we see that when $L_{y}\gg c/\omega_{p}$, the plasmoids become nearly insignificant compared to the large laminar exhausts that form. While plasmoids do perturb these exhausts (as noted in M. E. Innocenti et al. (2015)), they are too few in number to disrupt them: once the laminar exhausts stabilize, they persist for as long as we are able to simulate. Reconnection is still fast with the reconnection rate $R\equiv v_{\text{in}}/v_{A}\approx 0.10$, the same as for Run IV, and only somewhat lower than $R\approx 0.15$ for $\sigma=10$ (Run III).
@@
Thus, for low magnetization, the vast majority of inflowing particles avoid plasmoids altogether and simply proceed directly from the cold inflow to the hot exhaust. At this interface slow-mode shocks form, as can be confirmed via comparison with the six Rankine-Hugoniot jump conditions representing continuity in mass, normal momentum, energy, normal magnetic field, transverse momentum, and tangential electric field, respectively (N. K. Walia et al., 2024):
@@
Here $\rho$ represents the mass density, $\mathbf{v}$ the flow velocity, $P_{g}$ the gas pressure, and $\gamma=5/3$ the adiabatic index. The components normal and tangential to the shock are denoted by $n$ and $t$, respectively. Figure 3 shows these quantities along a horizontal cut across the exhaust boundary, far from the X-point. As one can see, these jump conditions are well satisfied in our simulations, in general agreement with the Petschek model (see N. K. Walia et al. (2022) for a similar analysis of a hybrid simulation). Note, however, that, contrary to H. E. Petschek (1964), these slow-mode shocks are not also switch-off shocks, consistent with earlier work (Y.-H. Liu et al., 2012). This is because some current is spread throughout the exhaust rather than being localized only at the shock front (Fig 3), resulting in the field lines gradually curving in the exhaust rather than bending sharply (Figure 2). Across these slow-mode shocks, the magnetic pressure is converted into particle pressure and the density rises to $n/n_{0}\approx 3$. Mass conservation allows us to determine the approximate outflow angle $\theta$, as a function of this ratio and the reconnection rate $R\approx 0.1$:
@@
This results in an outflow angle of $\theta\approx 4^{\circ}$, consistent with the simulation (Figure 2). While the relativistic case (Fig. 1a) does not have this sort of persistent exhaust, one can note smaller “minijets” trailing the plasmoids as they advect out (J. Ortuño-Macías & K. Nalewajko, 2020). The minijets display a drop in tangential B field across their boundary similar to a slow-shock transition.
@@ 3.2 Single to Multiple X-point Transition
In sharp contrast to this Petschek configuration, all of our $\sigma=10$ simulations (runs I and III), as well as those in larger 2D domains (e.g., L. Sironi et al., 2016), show large plasmoid chains with no significant exhausts. This sharp discrepancy is caused by the interplay between two competing processes operative in the current sheets. The first is the natural tendency for a reconnecting system to form wide-angle exhausts to ensure flow continuity (as in Equation (10)). In these exhausts, the incoming flow is heated and redirected to the outflow direction. The second is the fracturing of the thin current sheets in the diffusion regions into plasmoids as part of the secondary plasmoid tearing instability (D. A. Uzdensky et al., 2010; L. Sironi et al., 2016). These plasmoids, if sufficiently large and numerous, can disrupt the exhausts by physically displacing them. Conversely, the exhausts can suppress plasmoid formation because they widen the current sheet and thereby prevent further tearing.
To see why the $\sigma=10$ case is plasmoid-dominated and the $\sigma=0.3$ case is exhaust-dominated, let us consider spacetime diagrams showing the birth, growth, and outflow of plasmoids in the current sheet (Figure 4). The key difference between the two regimes is in plasmoid mergers, the process by which two plasmoids collide and form a larger plasmoid (for extensive analysis, see L. Sironi et al., 2016). In the $\sigma=10$ case (Figure 4(a)) this occurs constantly, with each plasmoid undergoing many mergers within its lifetime. In contrast, in the $\sigma=0.3$ case (Figure 4(b)) plasmoids seldom collide—the vast majority of plasmoids are simply born and then advect out of the diffusion region. Without mergers, these plasmoids never become very large—they simply grow modestly and then leave at the Alfvén speed (see also Figure 2). While the steady stream of numerous large plasmoids in the $\sigma=10$ case swiftly disrupts any nascent exhausts, the trickle of small plasmoids in the $\sigma=0.3$ is sufficiently feeble that exhausts can form unimpeded.
@@
Using the measured values of $\mathcal{F}$ and $w_{cs}$, we find $\mathcal{P}=0.10$ for Run I ($\sigma=10$) and $\mathcal{P}=0.05$ for Run II ($\sigma=0.3$). We can thus conclude that $\mathcal{P}\gtrsim 0.10$ corresponds to multiple X-point reconnection and $\mathcal{P}\lesssim 0.05$ to single X-point, separated by a critical transition parameter $\mathcal{P}_{c}\sim 0.075$. As $\mathcal{P}$ increases from $0.05$ to $0.10$ the stream of plasmoids intensifies as more mergers occur until any potential exhaust is rapidly dispersed. Since this expression is not unique to the particular system properties, we expect it to hold in electron-ion plasmas as well where similar current sheet structures are indeed observed (e.g., M. E. Rowan et al., 2017).
@@
To express this as a function of the magnetization, we make the following assumptions supported by our simulations. Firstly, we assume that the current sheet width $w_{cs}$ and density $n$ are independent of $\sigma$. This implies that $J^{2}\propto B_{0}^{2}\propto\sigma$. Secondly, we take $\langle\gamma\rangle\approx 1+\sigma/4$, the injection energy identified by S. R. Totorica et al. (2023). This results in the acceleration scaling as
@@ 4.1 Total Spectra
The electron energy spectra—specifically, the electron kinetic energy normalized to the magnetization—are compared in Figure 5. In the $\sigma=10$ case (e.g., L. Sironi & A. Spitkovsky, 2014; F. Guo et al., 2014), the plasmoid-dominated reconnection results in the establishment of a broad, hard power law (that is, at energies greater than those of the cold inflow component seen on the left of Figure 5). For the $\sigma=0.3$ case, we see two distinct components in the energized part corresponding to particles within the current sheet. First, on the left, we see a shoulder comprised of a flat spectrum followed by a rapid drop-off. At the end we see a small high-energy tail containing the most energetic particles which branches off from the shoulder at $\gamma\approx 1.7$. These features are caused by distinct particle acceleration mechanisms: the first “shoulder” component consists of particles that are energized in the exhaust and the second, more energetic, component consists of particles confined in plasmoids.
@@
In relativistic reconnection, particle acceleration is fundamentally stochastic—particles can transition from direct X-point acceleration to Fermi acceleration within plasmoids and back again. This results in the development of a smooth power-law spectrum. By contrast, in the $\sigma=0.3$ case, the regular, laminar structure of the reconnection region results in two distinct acceleration pathways which depend on where the particle first encounters the current sheet. The first category of particles follows the field lines into the diffusion region. There they are accelerated by the non-ideal electric fields (N. Bessho & A. Bhattacharjee, 2012) before being confined and energized by a compressing plasmoid (H. Hakobyan et al., 2021). The much larger second category of particles, instead of first encountering the current sheet at the X-point, does so at the slow-mode shocks at the boundary of the exhaust outflow. There they are pickup-accelerated and brought up to speed with the Alfvénic outflow, resulting in a thermal distribution. It is this latter pathway which ultimately dominates, especially when one considers realistic system sizes where $L_{y}\gg c/\omega_{p}$.

### v3

@@ 1 Introduction
Reconnection is a ubiquitous plasma process that rapidly converts magnetic energy into particle energy via rearrangement of the magnetic field topology. Originally posited to explain solar flares, reconnection now plays a key role in the theories of a multitude of astrophysical systems, from the heliosphere to compact objects. Observations and simulations suggest that reconnection can energize particles far beyond thermal energies, generating hard power-law energy spectra (sironi2014; guo2014). The conditions under which this particle acceleration occurs, the precise mechanisms responsible, and the shape of the resultant spectra have been intensively studied in recent years with the aid of particle-in-cell (PIC) simulations (li2021; guo2024; sironi2025). These simulations have enabled the study of kinetic reconnecting plasmas on increasingly large spatial and temporal scales.
@@
One of the key determinants of particle acceleration is the current sheet geometry: each particle acceleration mechanism requires certain configurations of electric and magnetic fields to operate. There are two broad categories of reconnecting current sheet geometries: single X-point and multiple X-point111also referred to as X-lines, the three-dimensional analogues of X-points. (ji2011). In the former, there is just one diffusion region where the field lines break and reconnect, while most of the reconnecting plasma proceeds directly from the inflow to the outflow exhaust. In the latter, the current sheet fragments due to tearing or plasmoid instability, resulting in a dynamic chain of magnetic islands with X-points in between (nuno2007). [^2] In this paper we shall use X-point and diffusion region interchangeably to refer to areas where non-ideal electric fields dissipate inflowing oppositely-aligned magnetic flux. Recent work has established the importance of direct electric field acceleration at X-points in relativistic plasmas as an injection mechanism for subsequent acceleration (sironi2022; totorica2023; gupta2025), a pathway which is more accessible in systems with more X-points. Furthermore, plasmoid-based Fermi acceleration mechanisms at later stages of energization (drake2006; guo2019; hakobyan2021) are also primarily operative in the multiple X-point regime. In contrast, pickup acceleration occurs mainly in the single X-point regime for particles that do not pass through the X-point (drake2009).
@@
Given the strong dependence of particle acceleration on the current sheet geometry, it would be useful to have a framework to predict the transition from single X-point reconnection to multiple X-point reconnection. In the MHD model, this is accomplished by the Lundquist number $S\equiv Lv_{A}/(\eta c/4\pi)$, where $L$ is the half-length of the current sheet, $v_{A}$ is the Alfvén speed, and $\eta$ is the resistivity, assumed to be constant. When $S>S_{c}\sim 10^{4}$, the reconnecting current sheet fragments, resulting in fast plasmoid-mediated multiple X-point reconnection (ji2011). Although recent strides have been made in understanding the tearing properties of collisionless current sheets (hoshino2020; schoeffler2024; akutagawa2025), there is no comparable metric to a priori predict the global reconnection geometry in collisionless systems, suggesting the need for kinetic studies.
@@
These questions also tie in closely to debates concerning the origin of fast reconnection. petschek1964 proposed that the slow Sweet-Parker reconnection (parker1957) could be made fast by flanking a small diffusion region with four slow-mode shocks responsible for the majority of energy dissipation. His model has proven controversial since, in the absence of additional physics to localize the diffusion region, the Petschek solution is unstable and devolves to that of Sweet and Parker in MHD simulations (uzdensky2000; forbes2013). On the other hand, slow-mode shocks have been observed in the Earth’s magnetotail for decades (e.g., feldman1987; saito1995; walia2024), providing observational evidence in favor of Petschek’s theory. Large-domain kinetic simulations would thus be well suited to address whether Petschek-type outflows can form, and, if so, under what conditions.
@@
where $B_{0}$ and $n_{0}$ are the magnetic field and number density of the upstream plasma, respectively. The magnetization represents the characteristic Lorentz factor that particles would gain if all the magnetic energy were dissipated into cold particles. Thus, $\sigma\gg 1$ indicates relativistic reconnection and $\sigma<1$ the subrelativistic regime. High-$\sigma$ systems are characterized by rapid plasmoid production, growth, and mergers (sironi2016). Since the rate of plasmoid formation increases with increasing $\sigma$, by lowering $\sigma$ below 1 we can address systems with lower tearing rates than those typically studied for pair plasmas. The result, as we shall demonstrate, is that while high-$\sigma$ simulations always yield multiple X-point plasmoid chains, those with low-$\sigma$ instead have just a single X-point. By comparing otherwise identical simulations that differ only in $\sigma$, we can understand the bifurcation of the system into single or multiple X-point regimes and study how this bifurcation affects the resultant particle acceleration.
In this investigation, we focus on two magnetization values: $\sigma=0.3$ and $\sigma=10$, which have (upstream) Alfvén speeds $v_{A}=c\sqrt{\sigma/(\sigma+1)}$ of $\approx 0.48c$ and $\approx 0.95c$, respectively. The $\sigma=10$ case is well within the relativistic reconnection regime and, as we shall see, the resulting reconnection rate, current sheet structure, and particle energy spectra are consistent with the existing literature (e.g., sironi2016; guo2024). This acts as a control with which we may contrast the $\sigma=0.3$ case. Although we focus on only these two cases, we expect that the key results for $\sigma=10$ also apply to $\sigma>10$, and the results of $\sigma=0.3$ apply to $\sigma<0.3$, with the range $0.3<\sigma<10$ presenting an intermediate regime.
@@ 2 Simulation Setup
We perform 2D simulations of reconnection using the electromagnetic PIC code TRISTAN-V2 (tristan). For all simulations, the mass ratio $m_{i}/m_{e}=1$, and there are 16 total particles per cell. The speed of light is set to $c=0.45$ cells per time step to satisfy the CFL condition. We vary the magnetization, skin depth, and domain size as described in Table 1.
@@ 3.1 Petschek-like Exhausts
Reducing the magnetization from $\sigma=10$ to $\sigma=0.3$ results in a stark change in the structure of the current sheet, as can be seen in Figure 1. In the $\sigma=10$ case, as expected, the tearing instability results in copious plasmoid formation, frequent plasmoid mergers, and rapid plasmoid growth. The result is a self-similar current sheet made up of plasmoids of various sizes, as described in detail in sironi2016. This is a clear case of multiple X-point reconnection wherein field lines reconnect at many locations along the current sheet. However, a primary X-point can be identified at $y\approx 360\;c/\omega_{p}$ during this timestep.
@@
The exhausts are even more pronounced in Run V that has a domain size ten times longer in the $y$-direction, shown in Figure 2. Here we see that when $L_{y}\gg c/\omega_{p}$, the plasmoids become nearly insignificant compared to the large laminar exhausts that form. While plasmoids do perturb these exhausts (as noted in innocenti2015), they are too few in number to disrupt them: once the laminar exhausts stabilize, they persist for as long as we are able to simulate. Reconnection is still fast with the reconnection rate $R\equiv v_{\text{in}}/v_{A}\approx 0.10$, the same as for Run IV, and only somewhat lower than $R\approx 0.15$ for $\sigma=10$ (Run III).
@@
Thus, for low magnetization, the vast majority of inflowing particles avoid plasmoids altogether and simply proceed directly from the cold inflow to the hot exhaust. At this interface slow-mode shocks form, as can be confirmed via comparison with the six Rankine-Hugoniot jump conditions representing continuity in mass, normal momentum, energy, normal magnetic field, transverse momentum, and tangential electric field, respectively (walia2024):
@@
Here $\rho$ represents the mass density, $\mathbf{v}$ the flow velocity, $P_{g}$ the gas pressure, and $\gamma=5/3$ the adiabatic index. The components normal and tangential to the shock are denoted by $n$ and $t$, respectively. Figure 3 shows these quantities along a horizontal cut across the exhaust boundary, far from the X-point. As one can see, these jump conditions are well satisfied in our simulations, in general agreement with the Petschek model (see walia2022 for a similar analysis of a hybrid simulation). Note, however, that, contrary to petschek1964, these slow-mode shocks are not also switch-off shocks, consistent with earlier work (liu2012). This is because some current is spread throughout the exhaust rather than being localized only at the shock front (Fig 3), resulting in the field lines gradually curving in the exhaust rather than bending sharply (Figure 2). Across these slow-mode shocks, the magnetic pressure is converted into particle pressure and the density rises to $n/n_{0}\approx 3$. Mass conservation allows us to determine the approximate outflow angle $\theta$, as a function of this ratio and the reconnection rate $R\approx 0.1$:
@@
This results in an outflow angle of $\theta\approx 4^{\circ}$, consistent with the simulation (Figure 2). While the relativistic case (Fig. 1a) does not have this sort of persistent exhaust, one can note smaller “minijets” trailing the plasmoids as they advect out (macias2020). The minijets display a drop in tangential B field across their boundary similar to a slow-shock transition.
@@ 3.2 Single to Multiple X-point Transition
In sharp contrast to this Petschek configuration, all of our $\sigma=10$ simulations (runs I and III), as well as those in larger 2D domains (e.g., sironi2016), show large plasmoid chains with no significant exhausts. This sharp discrepancy is caused by the interplay between two competing processes operative in the current sheets. The first is the natural tendency for a reconnecting system to form wide-angle exhausts to ensure flow continuity (as in Equation (10)). In these exhausts, the incoming flow is heated and redirected to the outflow direction. The second is the fracturing of the thin current sheets in the diffusion regions into plasmoids as part of the secondary plasmoid tearing instability (uzdensky2010; sironi2016). These plasmoids, if sufficiently large and numerous, can disrupt the exhausts by physically displacing them. Conversely, the exhausts can suppress plasmoid formation because they widen the current sheet and thereby prevent further tearing.
To see why the $\sigma=10$ case is plasmoid-dominated and the $\sigma=0.3$ case is exhaust-dominated, let us consider spacetime diagrams showing the birth, growth, and outflow of plasmoids in the current sheet (Figure 4). The key difference between the two regimes is in plasmoid mergers, the process by which two plasmoids collide and form a larger plasmoid (for extensive analysis, see sironi2016). In the $\sigma=10$ case (Figure 4(a)) this occurs constantly, with each plasmoid undergoing many mergers within its lifetime. In contrast, in the $\sigma=0.3$ case (Figure 4(b)) plasmoids seldom collide—the vast majority of plasmoids are simply born and then advect out of the diffusion region. Without mergers, these plasmoids never become very large—they simply grow modestly and then leave at the Alfvén speed (see also Figure 2). While the steady stream of numerous large plasmoids in the $\sigma=10$ case swiftly disrupts any nascent exhausts, the trickle of small plasmoids in the $\sigma=0.3$ is sufficiently feeble that exhausts can form unimpeded.
@@
Using the measured values of $\mathcal{F}$ and $w_{cs}$, we find $\mathcal{P}=0.10$ for Run I ($\sigma=10$) and $\mathcal{P}=0.05$ for Run II ($\sigma=0.3$). We can thus conclude that $\mathcal{P}\gtrsim 0.10$ corresponds to multiple X-point reconnection and $\mathcal{P}\lesssim 0.05$ to single X-point, separated by a critical transition parameter $\mathcal{P}_{c}\sim 0.075$. As $\mathcal{P}$ increases from $0.05$ to $0.10$ the stream of plasmoids intensifies as more mergers occur until any potential exhaust is rapidly dispersed. Since this expression is not unique to the particular system properties, we expect it to hold in electron-ion plasmas as well where similar current sheet structures are indeed observed (e.g., rowan2017).
@@
To express this as a function of the magnetization, we make the following assumptions supported by our simulations. Firstly, we assume that the current sheet width $w_{cs}$ and density $n$ are independent of $\sigma$. This implies that $J^{2}\propto B_{0}^{2}\propto\sigma$. Secondly, we take $\langle\gamma\rangle\approx 1+\sigma/4$, the injection energy identified by totorica2023. This results in the acceleration scaling as
@@ 4.1 Total Spectra
The electron energy spectra—specifically, the electron kinetic energy normalized to the magnetization—are compared in Figure 5. In the $\sigma=10$ case (e.g., sironi2014; guo2014), the plasmoid-dominated reconnection results in the establishment of a broad, hard power law (that is, at energies greater than those of the cold inflow component seen on the left of Figure 5). For the $\sigma=0.3$ case, we see two distinct components in the energized part corresponding to particles within the current sheet. First, on the left, we see a shoulder comprised of a flat spectrum followed by a rapid drop-off. At the end we see a small high-energy tail containing the most energetic particles which branches off from the shoulder at $\gamma\approx 1.7$. These features are caused by distinct particle acceleration mechanisms: the first “shoulder” component consists of particles that are energized in the exhaust and the second, more energetic, component consists of particles confined in plasmoids.
@@
In relativistic reconnection, particle acceleration is fundamentally stochastic—particles can transition from direct X-point acceleration to Fermi acceleration within plasmoids and back again. This results in the development of a smooth power-law spectrum. By contrast, in the $\sigma=0.3$ case, the regular, laminar structure of the reconnection region results in two distinct acceleration pathways which depend on where the particle first encounters the current sheet. The first category of particles follows the field lines into the diffusion region. There they are accelerated by the non-ideal electric fields (bessho2012) before being confined and energized by a compressing plasmoid (hakobyan2021). The much larger second category of particles, instead of first encountering the current sheet at the X-point, does so at the slow-mode shocks at the boundary of the exhaust outflow. There they are pickup-accelerated and brought up to speed with the Alfvénic outflow, resulting in a thermal distribution. It is this latter pathway which ultimately dominates, especially when one considers realistic system sizes where $L_{y}\gg c/\omega_{p}$.